OpenStudy (loser66):

Confuse!! Please, help Problem: Where is the function differentiable? \(x^2+y^2 +2ixy\) Where is it analytic?

OpenStudy (loser66):

\(U_x = 2x = V_y \\U_y = 2y = -V_x = -2y \iff y =0\) Hence the function is differentiable on x axis. My question: is it not that it is analytic on x axis also? why the answer is : "No where" ?

OpenStudy (anonymous):

hmm you sure this is right?

OpenStudy (loser66):

What is your suspicion?

OpenStudy (anonymous):

checking the partials although it can't be that hard, i think it must be of by a minus sign somewhere

OpenStudy (loser66):

\(U(x,y) = x^2 +y^2 \\V(x,y) = 2xy\) \(U_x = 2x\\U_y= 2y\\V_x = 2y\\V_y = 2x\) I am sure about that.

OpenStudy (anonymous):

then you are off by a minus sign for cauchy riemman right?

OpenStudy (loser66):

\(U_x = V_y=2x\\U_y = -V_x\iff 2y = -2y \iff y =0\) I am sure about that also.

OpenStudy (loser66):

Yes,

OpenStudy (anonymous):

for cauchy riemann \(U_y=-V_x\)

OpenStudy (loser66):

Yes

OpenStudy (anonymous):

but is isn't

OpenStudy (loser66):

Hence, the conclusion is the function is differentiable when y =0

OpenStudy (loser66):

That is it is differentiable on Real axis, right?

OpenStudy (anonymous):

yeah on the real axis it is \(f(x)=x^2\)

OpenStudy (loser66):

But for part b) Where is it analytic, From above it is differentiable on Real, it should be analytic on Real also, but the answer is "No where"

OpenStudy (loser66):

It is not f (z)

OpenStudy (anonymous):

differentiable means the derivative exists from any direction

OpenStudy (loser66):

f(z) = x^2 +y^2 +2ixy,

OpenStudy (loser66):

and the whole thing differentiable at the point on Real axis only. We can't take off y from f, right?

OpenStudy (anonymous):

no just like on the real line differentiable means the limit exists from the left and the right, in complex plane the limit must exist from any direction

OpenStudy (anonymous):

btw if they didn't mention it, as i remember correctly "analytic" means differentiable, but also means you can write the function as a function of \(z\)

OpenStudy (loser66):